### v1

# KCL-PH-TH/2025-19

Abstract
We investigate two classes of dark matter (DM) candidates, sub-GeV particles and primordial black holes (PBHs), that can inject low-energy electrons and positrons into the Milky Way and leave observable signatures in the X-ray sky. In the case of sub-GeV DM, annihilation or decay into $e^{+}e^{-}$ contributes to the diffuse sea of cosmic-ray (CR) leptons, which can generate bremsstrahlung and inverse Compton (IC) emission on Galactic photon fields, producing a broad spectrum from X-rays to $\gamma$-rays detectable by instruments such as eROSITA and XMM-Newton. For PBHs with masses below $\sim 10^{17}$ g, Hawking evaporation similarly yields low-energy $e^{\pm}$, leading to comparable diffuse emission. Using the first data release from eROSITA and incorporating up-to-date CR propagation and diffusion parameters, we derive new constraints on both scenarios. For sub-GeV DM, we exclude thermally averaged annihilation cross sections in the range $\sim 10^{-27}$$10^{-25}\ \mathrm{cm^{3}/s}$ and decay lifetimes of $\sim 10^{24}$$10^{25}$ s for masses between 1 MeV and 1 GeV, with eROSITA outperforming previous X-ray constraints below $\sim 30$ MeV. For asteroid-mass PBHs, we set new bounds on the DM fraction based on their Hawking-induced emission. Finally, we revisit earlier constraints from XMM-Newton, finding that they were approximately four orders of magnitude too stringent due to the use of the instrument’s geometric solid angle rather than its exposure-weighted solid angle. Upon using the exposure-weighted solid angle, we show that the revised XMM-Newton limits are slightly weaker than those from eROSITA.
1 Introduction
The nature of dark matter (DM) remains one of the most compelling open problems in modern physics. A wealth of astrophysical and cosmological observations indicate that approximately 85% of the matter content of the Universe is composed of DM [1, 2]. Yet, its fundamental properties continue to elude direct detection. While weakly interacting massive particles (WIMPs) have long been leading candidates, their persistent non-observation [2] has intensified the exploration of alternative scenarios, including sub-GeV DM [3, 4, 5, 6] and primordial black holes (PBHs) [7, 8, 9].
Sub-GeV DM candidates, with masses in the MeV–GeV range, are theoretically well-motivated and arise in a variety of extensions of the Standard Model [5, 10, 11]. However, their low mass poses detection challenges. Direct detection experiments lose sensitivity in this regime, while indirect searches are hindered by charged particle suppression in the heliosphere and the limited sensitivity of existing $\gamma$-ray instruments in the keV–GeV band, the so-called “MeV gap” [12, 10, 13]. Nevertheless, DM annihilation or decay into $e^{\pm}$ can inject low-energy electrons and positrons into the interstellar medium. These secondaries interact with Galactic radiation fields via bremsstrahlung radiation and inverse Compton (IC) scattering, producing X-rays in the keV to MeV range [14], which offers a valuable observational window.
PBHs, originally proposed decades ago [15, 16], have reemerged as viable DM candidates, particularly in light of null results from particle-based searches. PBHs with masses $M_{\rm BH}\lesssim 10^{17}$ g can emit high-energy particles via Hawking evaporation [17, 18, 8]. Those below $10^{16}$ g can emit charged particles that generate a diffuse X-ray signal, similarly to sub-GeV DM scenarios. While these low-energy particles are difficult to detect directly on Earth due to solar modulation, observations from Voyager 1 beyond the heliopause have enabled constraints to be derived [19, 20].
X-ray observations thus offer a uniquely complementary approach to probing both sub-GeV DM and PBHs. They are particularly sensitive to models involving low-energy $e^{\pm}$ injection and probe the MeV gap inaccessible to most $\gamma$-ray telescopes. Past studies using instruments like XMM-Newton have searched for diffuse X-ray signals from light DM [14] and PBHs [21]. Despite the long exposure times achievable with pointing telescopes, their narrow fields of view limit coverage to specific sky regions, constraining their ability to study large-scale emission or high-latitude backgrounds.
The Extended ROentgen Survey with an Imaging Telescope Array (eROSITA) represents a major advance for such searches. Launched aboard the Spectrum-Roentgen-Gamma (SRG) satellite in 2019 [22], eROSITA provides the most sensitive full-sky coverage in the 0.2–8 keV range, including the first complete view of the X-ray sky above 2 keV [23]. Even with only half its planned surveys completed, eROSITA already surpasses ROSAT in soft X-ray sensitivity, owing to its order-of-magnitude larger effective area above 0.3 keV and its degree-scale field of view. These features, including large sky coverage, high angular and spectral resolution, and deep exposure, make eROSITA ideally suited to search for both secondary X-ray emission from DM-induced $e^{\pm}$, and for Hawking radiation signatures from PBH evaporation across the Galactic halo.
While the Galactic Centre (GC) offers enhanced DM densities and signal brightness, eROSITA’s full-sky reach allows constraints to be derived from high-latitude regions where astrophysical backgrounds are lower. Prior eROSITA analyses have targeted sterile neutrinos and axion-like particles using the eFEDS field [24, 25, 26], but no comprehensive study has yet leveraged eROSITA data to constrain sub-GeV DM or PBHs with full consideration of CR propagation and secondary emission.
In this work, we address this gap by placing new constraints on light DM and sub-asteroid mass PBHs using eROSITA’s full-sky X-ray maps. We model continuum emission from DM-induced low-energy $e^{\pm}$ and Hawking evaporation from PBHs, deriving upper limits on the annihilation cross section, decay lifetime, and PBH DM fraction across a wide mass range. We then compare these constraints with existing bounds from other probes, demonstrating the complementarity and reach afforded by eROSITA. Additionally, we revisit constraints derived from XMM-Newton data [27], showing that applying the geometric solid angle instead of the exposure-weighted solid angle led to flux differences of approximately four orders of magnitude. After addressing this, the XMM-Newton limits are found to be weaker than those from eROSITA, realigning the landscape of X-ray DM constraints.
This paper is structured as follows. In Section 2, we describe the eROSITA dataset. Section 3 details our signal modelling and data analysis framework. In Section 4, we present our results and compare them with existing constraints. Section 5 discusses the reanalysis of XMM-Newton data and the impact of correcting the exposure estimate. Finally, we summarise our findings and conclude in Section 6.
2 eROSITA diffuse emission - Data analysis
Table 1: Comparison of eROSITA with other X-ray satellites covering the keV band.
Telescope Energy range Effective area FoV Spatial res. Sensitivity
(keV) (cm2 at 1 keV) (°) (arcsec) (erg cm-2 s-1)
eROSITA 0.2-10 1237 1 $<$10 $\sim$5$\times$10-14
(eRASS1, $<$2.3 keV)
XMM-Newton 0.15-12 $\sim$ 1200 (250) 0.5 6 $\sim$10-14
total (single MOS) (after 10 ks)
Chandra 0.1-10 555 0.5 0.5 4$\times$10-15
(after 100 ks)
ROSAT 0.1-2.4 400 2 $\sim$30 1.5$\times$10-13
This study is motivated by the technological advancements introduced by the recently launched eROSITA telescope in the X-ray domain. eROSITA comprises seven parallel-aligned X-ray telescope modules equipped with single chip-PN CCD detectors, which are similar to but improved upon those of XMM-Newton PN (discussed below). Each module features a $1\degree$ field of view (FoV) and individually hosts 54 nested mirror shells [23]. This configuration offers our most sensitive view of the X-ray sky.
Although eROSITA is not a pointed X-ray instrument and therefore does not broadly permit X-ray observations lasting tens or hundreds of kiloseconds, it is significant to demonstrate its major advancements compared to other X-ray instruments operating in the same energy range (or there is at least a strong overlap) of $\sim 0.2$-$10$ keV, such as ROSAT (the predecessor of the eROSITA all-sky survey), XMM-Newton and Chandra.
The combined effective area from the seven eROSITA telescope modules is substantially greater than that of ROSAT and Chandra, though only slightly larger compared to XMM-Newton (MOS+PN). However, what distinguishes eROSITA is its enhanced spectral resolution compared to XMM-Newton PN and the combination of the large effective area with the $1\degree$ FoV, which introduces the so-called grasp concept (the product of the effective area and FoV), which is significantly higher (by many times) than that of any current X-ray survey in the 0.3-3.5 keV energy range, as illustrated in Ref. [23] (Fig. 10). For a direct comparison, we summarise the aforementioned instruments’ capabilities in Tab. 1.
Although the initial objective was to conduct eight all-sky surveys over a four-year period, only four surveys have been successfully completed to date. The corresponding datasets are not publicly accessible in their entirety. As such, this research utilises the newly available eRASS1 X-ray all-sky survey datasets, which became publicly available in February 2024, and are characterised by unprecedented sensitivity. The aim is to exploit the soft broadband eROSITA maps (0.3-2.3 keV) for the purposes of this DM search.
The calibrated eRASS1 data were processed using the eROSITA Standard Analysis Software (eSASS) [28], adhering to all standard recommendations for data reduction from the eROSITA collaboration, which include exposure correction and the exclusion of bad time intervals and flaring events. Since the eROSITA survey consists of 4700 overlapping distinct sky tiles, of $3.6\degree\times 3.6\degree$ each, the map reported in this study was produced by assembling the exposure corrected count images of the adjacent sky tiles that cover the selected region to be studied.
The maps are produced in HEALPix format, using the healpy package [29], with a pixel size of $0.22\degree\times 0.22\degree$ in units cts/s/$\mathrm{deg}^{2}$ (and subsequently converted to keV/sr/$\mathrm{cm^{2}}$/s for the purposes of the analysis) after properly filtering out point sources (according to the eRASS1 source catalog [30]) above a selected flux threshold (point sources with ML_FLUX_P2+ML_FL UX_P3 in the eRASS1 source catalog above $10^{-13}$ were removed) and removing the particle background. This threshold is selected such that 19% of the cosmic X-ray background is removed, as discussed in Fig. 12 of the main data release collaboration paper [30]. Overall, the individual steps outlined by the eROSITA collaboration in Section 2 of Ref. [31] were adapted to obtain the corresponding diffuse X-ray map from the selected region.
From this analysis, we obtain a full-sky map of the unresolved-diffuse emission observed by eROSITA in the $\sim 0.1$-$2.3$ keV band. Given the angular resolution of the telescope and its poor sensitivity for source detection (poor, at least, in comparison to XMM-Newton or Chandra), this emission results from the sum of the truly diffuse emission, originated from the IC emission of CR electrons off the interstellar radiation fields (ISRFs) [32] and the unresolved sources below the detection threshold of eROSITA. Therefore, with more observation time it will be possible to lower the diffuse background observed by eROSITA and improve current constraints.
Subsequently, we divide the Galaxy into 30 sub-regions that consist of $6^{\circ}$ wide concentric rings around the GC, following the strategy followed by Ref. [27], with the XMM-Newton data. This allows a consistent comparison between the eROSITA and the XMM-Newton diffuse emission in the $0.1$ to $\sim 10$ keV range, that already reveals that previous XMM-Newton constraints have been overestimated. However, since the X-ray emission absorption by interstellar gas becomes more important at lower energy, we ignore the region of latitude $|b|<5^{\circ}$ to mask the Galactic plane, as opposed to the analysis XMM-Newton data, which used $|b|<2^{\circ}$. Therefore, our result leads to more conservative constraints. We show an example of Mollewide projections of the eROSITA HEALPix data and the rings used in Fig. 1.
Figure 1: Mollweide projections of the X-ray sky observed by eROSITA. The left panel shows the map with the Galactic plane mask applied ($|b|<5^{\circ}$), while the right panel displays the concentric $6^{\circ}$ ringed regions used for subdivision of the Galaxy that will be used to derive constraints on dark matter, indicating the observed count rate as a color bar.
3 Continuum signals from dark matter
3.1 Sub-GeV dark matter annihilation and decay
In this work we implement the injection and propagation of $e^{\pm}$ from light DM, using a modified version of the DRAGON2 code [33, 34], which solves the CR diffusion-advection-loss equation [35]. The DM injection spectra are taken from Ref. [10], where we refer the reader for a comprehensive discussion of their production spectra. For annihilation or decay, the crucial input of interest is the injection rate of electrons and positrons per unit volume, which is denoted $Q_{e}(\vec{x},E)$ and takes the form
$Q_{e}=\begin{cases}\frac{\langle\sigma v\rangle}{2}\left(\frac{\rho_{\chi}(\vec{x})}{m_{\chi}}\right)^{2}\frac{dN_{e}}{dE_{e}}&\text{(annihilation)}\\ \Gamma\left(\frac{\rho_{\chi}(\vec{x})}{m_{\chi}}\right)\frac{dN_{e}}{dE_{e}}&\text{(decay)},\end{cases}$ (3.1)
with $\langle\sigma v\rangle$, $\Gamma$, $\rho_{\chi}$ and $\frac{dN_{e}}{dE_{e}}$ denoting the annihilation cross section, decay rate, DM density profile and spectrum, respectively. The above formulae assumes Majorana DM particles, an additional factor of 1/2 has to be included for annihilation in the case of Dirac fermionic DM since $\langle\sigma v\rangle$ has to be averaged over particles and antiparticles. In this work, we will proceed assuming Majorana DM as our base case, which can easily be rescaled in the case of Dirac DM.
We restrict our analysis to the channel of direct $e^{+}e^{-}$ production, from a minimum kinematically allowable mass of $m_{e}$ ($2m_{e}$) for annihilation (decay) respectively. In both cases, we take the maximum mass of DM to be $1000$ MeV. We use a 2D axisymmetric Galactic model with a NFW DM profile [36] representing our base case. Other decay channels including the production of pions and muons can also be considered, but the constraints become much weaker than in previous analyses for the high DM masses needed. The propagation parameters used in this work are obtained from a fit to AMS-02 data [37, 38], and are listed in Table I of Ref. [32], where more of the technical details are given. The interplay between diffusion and energy losses is key at sub-GeV energies. Energy losses in the DRAGON2 code include ionisation, Coulomb, bremsstrahlung, IC, synchrotron and in-flight positron annihilation processes. The spatial and energy resolution employed is $\sim 150$ pc and $5$%.
The resulting $e^{\pm}$ distributions are passed to the HERMES code [39] to compute IC X-ray emission, integrating over the ISRFs [40]. Although IC losses have little effect on the $e^{\pm}$ spectra, the emission dominates the X-ray signal, especially below the MeV range that is relevant for experiments such as eROSITA and XMM-Newton. We estimate the ISRF modelling uncertainty to be below 30% [41], and we separately account for internal bremsstrahlung, also known as final-state radiation (FSR), as in Ref. [14]. Full simulation inputs and outputs are available in Ref. [42].
In Fig. 2 we show examples of the expected X-ray flux spectra in eROSITA in the ring 3 region of the Galaxy for DM masses $m_{\chi}$ of 1, 10, 100, 1000 $\mathrm{MeV}$ for both annihilation and decay, with fixed $\langle\sigma v\rangle=2.4\times 10^{-26}\ \mathrm{cm^{3}/s}$ and $\tau=1\times 10^{24}\ \mathrm{s}$ respectively (colored lines). We show the measured flux of eROSITA for comparison (black points).
On top of the IC emission, we also consider the contribution from FSR for every mass, which is computed following Ref. [14]. This contribution becomes more relevant for lower masses, being important only below $\sim 10$ MeV. In our case, this contribution remains negligible when taking into account reacceleration, as shown in Ref. [32].
Figure 2: Comparison of eROSITA data (black points) in Ring 3 to the predicted particle DM or PBH evaporation induced X-ray signal for annihilation with $\langle\sigma v\rangle=2.4\times 10^{-26}\ \mathrm{cm^{3}/s}$ (top left) and decay with $\tau=1\times 10^{24}\ \mathrm{s}$ (top right) and for PBHs with $f_{\mathrm{PBH}}=1$ (bottom). We show the predicted flux for different particle DM and PBH masses, as specified in the legends.
3.2 Primordial black hole evaporation
Low mass black holes are theoretically expected to evaporate via Hawking radiation, emitting particles with energies up to the black hole temperature $T$ [43, 44], which depends on its mass $M$, spin parameter $a\equiv J/M$ and charge
$T=\frac{1}{2\pi}\left(\frac{r_{+}-M}{r_{+}^{2}+a^{2}}\right),\quad r_{+}=M+\sqrt{M^{2}-a^{2}}.$ (3.2)
For Schwarzschild black holes ($a=0$), this reduces to $T=1/(8\pi M)$. The primary particle spectrum for species index $i$ is then
$\frac{d^{2}N_{i}}{dtdE_{i}}=\frac{1}{2\pi}\sum_{\text{d.o.f.}}\frac{\Gamma_{i}(E_{i},M,a^{\star})}{e^{E^{\prime}_{i}/T}\pm 1},$ (3.3)
where the energy $E^{\prime}_{i}\equiv E_{i}-m\Omega$, $\Omega=a^{\star}/(2r_{+})$, and $\Gamma_{i}$ are greybody factors accounting for gravitational redshift. We compute these spectra using BlackHawk v2.2 [45, 46].
To model sub-GeV $e^{\pm}$ relevant for X-ray production, we use Hazma [47] for secondary particle spectra in the domain $M\sim 10^{14.5}$$10^{17.5}$ g, where evaporation produces $e^{\pm}$, $\gamma$, and $\nu_{e,\mu,\tau}$. This accounts for decays of $\mu^{\pm}$, $\pi^{0,\pm}$ and FSR.
The injection rate of $e^{\pm}$ from PBHs per unit volume is
$Q_{e}(E_{e},\vec{x})=f_{\textrm{PBH}}\rho_{\textrm{DM}}(\vec{x})\int_{M_{\textrm{min}}}^{\infty}\frac{dM}{M}\frac{dN_{\textrm{PBH}}}{dM}\frac{d^{2}N_{e}}{dtdE_{e}},$ (3.4)
where $f_{\textrm{PBH}}$ is the PBH fraction of DM, and $M_{\textrm{min}}\approx 7.5\times 10^{14}$ g is the minimum mass of PBHs surviving today. For simplicity we adopt a monochromatic mass distribution ($\frac{dN_{\textrm{PBH}}}{dM}=\delta(M-M_{\textrm{PBH}})$) and assume non-rotating black holes, which constitute the most conservative scenario.
3.3 Dark matter density profiles
In this analysis, we investigate three widely studied density profiles for describing the distribution of DM: the Navarro-Frenk-White (NFW) profile [36, 48] as our base case, a contracted Navarro-Frenk-White (cNFW) profile for a more cuspy scenario, which accounts for baryonic contraction [49, 50] and the isothermal profile [51, 52] for a more cored conservative scenario. These three benchmarks are useful to understand the uncertainties in the constraints derived from the assumption of the DM density distribution.
The NFW profile, derived from numerical N-body simulations of cold DM halos, characterises the DM density as
$\rho_{\mathrm{NFW}}(r)=\frac{\rho_{s}}{\frac{r}{r_{s}}\left(1+\frac{r}{r_{s}}\right)^{2}},$ (3.5)
where $\rho_{s}$ is the characteristic density, set to from the assumption that the local DM density is of $0.4$ GeV/$\mathrm{cm^{3}}$ and $r_{s}=20~{}\text{kpc}$ [53].
However, baryonic processes, such as gas cooling and subsequent star formation, are expected to contract the DM halo inwards, leading to higher central densities. To account for this baryonic contraction, we adopt the cNFW profile. This modified profile uses the standard adiabatic contraction formalism introduced by Ref. [49] and refined by Ref. [50]. The cNFW profile can significantly enhance the central DM density, thus strengthening observational signatures and constraints. The contracted NFW profile can be written
$\rho_{\mathrm{cNFW}}(r)=\frac{\rho_{s}}{\left(\frac{r}{r_{s}}\right)^{\gamma}\left(1+\frac{r}{r_{s}}\right)^{3-\gamma}},$ (3.6)
where the normalisation $\rho_{s}$ is again given by fixing the local DM density to $0.4$ GeV/$\mathrm{cm^{3}}$, the scale radius is the same as for the NFW case and we set the inner slope to $\gamma=1.4$.
Additionally, we consider an isothermal profile as a pessimistic scenario. The isothermal DM profile is a spherically symmetric, cored density profile that assumes DM behaves like an ideal gas in thermal equilibrium, leading to a flat central density core. The expression for the density profile can be written
$\rho_{\textrm{iso}}(r)=\frac{\rho_{0}}{1+\left(\frac{r}{r_{c}}\right)^{2}},$ (3.7)
with a typical scale radius of $r_{c}=3.5$ kpc [33, 53]. This profile is often employed to explore scenarios with minimal astrophysical signals arising from DM annihilation and decay, thereby providing conservative limits on DM properties.
In subsequent sections, we consider the NFW profile as our base case, but also apply the cNFW and isothermal profiles when parameterising DM profile uncertainties, thereby evaluating the impact of uncertainty in the DM concentration in the centre of the Galaxy when evaluating DM indirect detection signals.
4 Constraints from eROSITA data
Following the strategy from previous works such as Refs. [14, 32], to derive constraints on the DM thermally averaged cross-section $\langle\sigma v\rangle$ and lifetime $\tau$ for annihilation and decay into $e^{+}e^{-}$ respectively, as well as the fractional abundance $f_{\mathrm{PBH}}=\rho_{\mathrm{PBH}}/\rho_{\mathrm{DM}}$ in the case of PBHs, we use a combined analysis of all rings by using the chi-squared statistic,
[…]
Figure 3: Comparison of the $95\%$ confidence bounds on annihilating (top left panel) and decaying (top right panel) DM with previous CR constraints. The bottom left panel shows constraints on the fraction of PBHs as DM derived in this work. Finally, in the bottom right panel, we show a comparison of the eROSITA decay constraints with other probes. In all the cases we report the results using the best-fit propagation parameters for eROSITA (orange line), as well as uncertainties arising from using pessimistic and optimistic propagation parameters (reddish band), alongside other existing constraints (green lines).
To account for uncertainties in propagation, we evaluate constraints under three different propagation scenarios: standard, optimistic, and pessimistic (following Refs. [32, 21]). These scenarios correspond to different values of key propagation parameters. Specifically, we vary the Alfvén velocity $v_{A}$ as 13.43, 40, and 0 km s-1, and the halo height $H$ as 8, 16, and 4 kpc, for the standard, optimistic, and pessimistic cases, respectively. Since the diffusion coefficient $D$ is constrained by the ratio $D/H$, it is adjusted accordingly to maintain consistency with the secondary-to-primary CR ratios measured by AMS-02 [54, 55, 38].
In Fig. 3, we show the constraints on the DM annihilation cross section, $\langle\sigma v\rangle$, and decay lifetime, $\tau$, into electrons and positrons, assuming a nominal NFW DM profile, using our combined ring analysis (orange line). The orange uncertainty bands represent those arising from varying of the propagation parameters in the scenarios described above. We also show existing constraints from INTEGRAL, NuSTAR and Suzaku (green lines) [14, 56, 57]. Compared to these probes, eROSITA provides the strongest constraints for DM masses up to a few tens of MeV. If we take constraints from rings individually, we find that rings 1 - 3 provides the strongest constraints, nearly identical to constraints derived from the full ring analysis for annihilation, as opposed to decay which benefits more from using all rings. As we observe, the uncertainties are larger for lower DM masses, which is due to the impact of reacceleration in the $e^{\pm}$ products: the lower their energy, the more prone they are to being boosted by the interaction with plasma waves [58, 59], as extensively reported and discussed in Refs. [32, 13].
To highlight the importance of these constraints, we show a comparison of the decay constraints obtained with eROSITA in comparison with those obtained from cosmological measurements, heating of dwarf galaxies and direct observations of CRs (in particular, from Voyager-1) and diffuse X- and $\gamma$-rays. We are not showing constraints from the 511 keV line [60] in this plot, which can reach lifetimes of $10^{29}$ s for $1$ MeV DM particles. The cosmological constraints are from CMB measurements and are obtained from the study of anisotropies in the CMB power spectrum [61, 62], the constraint from Leo-T dwarf galaxy was derived requiring the rate of heat injection by DM to not exceed the ultra-low radiative cooling rate of gas observed [63], and those from diffuse $\gamma$-rays are constraints from the FSR emission associated with DM [64] (obtained with COMPTEL, INTEGRAL and HEAO-I). Meanwhile, the constraint from Voyager-1 data is obtained by comparing the expected interstellar flux of electrons and positrons from DM with the flux measurements on board the probe. For the latter, it is important to note that in conservative conditions where reacceleration is very low, no limit can be obtained for DM masses below $\sim 20$ MeV. As we see, the eROSITA constraints are especially competitive below $10$ MeV, although at higher masses they are weaker. However, we emphasise that these limits are quite conservative, given that we directly mask the inner ±$\pm 5^{\circ}$ around the Galactic plane [^1] When doing the analysis adopting a ±$\pm 2^{\circ}$ mask, our constraints improve by no more than a factor of 2.. On top of this, we note that including the astrophysical background components could lead to up to even an order of magnitude stronger constraints from eROSITA.
Figure 4: Comparison of the $95\%$ confidence bounds on annihilating (left panel) and decaying (right panel) DM with bands indicating the uncertainty from the DM density. The central line shows constraints from choosing the standard NFW profile and assuming a local DM density of $0.4$ GeV/cm3. Shaded bands show uncertainty arising from considering a local DM density ranging from $0.3$ to $0.6$ GeV/cm3 (darker bands) and varying the density profile from cNFW to isothermal distributions (lighter bands).
We also quantify the uncertainty in the constraint arising from the DM density estimation. In particular, we show a band that represents the uncertainty due to the DM density profile adopted (Sec. 3.3) and another band that represents the uncertainty in determining the local DM density. To achieve this, we simulate the DM continuum emission for a contracted NFW (cNFW) profile with $\gamma=1.4$ which has a steeper profile than the standard NFW, and an isothermal distribution, which is more shallow. As discussed above, these represent two opposite and extremal scenarios that can significantly boost or suppress the diffuse X-ray emission generated by DM.
In these cases, we use the standard propagation parameters, as explained above. In Fig. 4, we show the DM constraint for NFW (orange line), and the uncertainty derived from the adoption of cNFW and isothermal distributions (orange band) for both annihilation and decay scenarios. In the case of annihilation, the DM induced flux goes as $\phi_{\mathrm{DM}}\sim\rho_{\mathrm{DM}}^{2}$, making the profile uncertainties more dominant. Meanwhile, for decay, the flux dependence is weaker, varying as $\phi_{\mathrm{DM}}\sim\rho_{\mathrm{DM}}$, and the uncertainty from the adopted profile is much lower, as expected. From this comparison we conclude that while above a few MeVs the DM density profile is the main source of uncertainty in these limits, below 10 MeV, propagation uncertainties dominate.
Finally, in the bottom panel of Fig. 3, we also report constraints on PBHs in the mass range $10^{14}\ \mathrm{g}\lesssim m_{\mathrm{PBH}}\lesssim 10^{17}\ \mathrm{g}$ from Hawking evaporation of charged particles (orange line). As commented above, these constraints are derived assuming monochromatic and non-rotating PBHs. The orange band again represent uncertainties arising from propagation. Here, we also compare our result from eROSITA with other leading constraints, including the 511 keV line, XMM-Newton and Voyager 1 (as green lines). In comparison to these probes, eROSITA diffuse background observations do not provide competitive constraints.
5 Revising XMM-Newton constraints
As noted in the previous section, the DM constraints derived from eROSITA data are significantly weaker than those reported in earlier analyses of XMM-Newton observations [14, 32]. At first glance, this appears counterintuitive. While eROSITA probes lower energy bands than XMM-Newton, where the latter suffers from increased instrumental noise, one would still expect comparable fluxes in the same regions of the sky when interpreted correctly.
We believe this discrepancy arises from a misinterpretation in previous analyses regarding the solid angle associated with the data presented in Ref. [27]. In particular, earlier works assumed the reported fluxes were distributed over the full geometric solid angle of each ring. However, the data in Ref. [27] are normalised to the exposure-weighted average solid angle $\Omega_{\mathrm{exp}}$, which accounts for the non-uniform exposure across the ring due to the telescope’s observational footprint [^2] A Jupyter notebook tutorial on how to interpret the data is now provided in the original Github repository.. This is defined
[…]
where $T_{\mathrm{eff}}$ is the exposure time in the angular direction and $T_{\mathrm{tot}}$ is the normalising total exposure time of the experiment, integrating over all angular directions in the ring region R. Crucially, these exposure-weighted solid angles are several orders of magnitude smaller than the corresponding geometric solid angles for each ring. For example, the geometric solid angle of ring 3 is $\Omega\approx 0.16\ \mathrm{sr}$, whereas the exposure-weighted solid angle for XMM-Newton in ring 3 is $\Omega_{\mathrm{exp}}\approx 3.5\times 10^{-5}\ \mathrm{sr}$.
As a result, when DM-induced X-ray fluxes were computed assuming the geometric solid angle, the flux was overestimated by several orders of magnitude relative to the flux actually reported by XMM-Newton, which properly accounts for the limited exposure. This can lead to spuriously strong DM constraints being derived.
To emphasise this, in Fig. 5 we plot the ring 3 fluxes (using a comparable $|b|<2^{\circ}$ inner mask of the Galactic plane) recorded for both eROSITA (orange points) and XMM-Newton assuming the exposure-weighted average solid angle (black points) as intended in the data analysis and the geometric solid angle (grey points) in the energy range 00-$6$ keV. We use experiment independent units of $\mathrm{keV\ s^{-1}\ sr^{-1}\ cm^{-2}}$ to produce compatible spectra. We see that using the geometric solid angle significantly underestimates the true flux and, in fact, XMM-Newton should provide somewhat weaker or comparable constraints to eROSITA once this has been addressed.
Figure 5: Comparison of ring 3 flux for eROSITA (orange points) and XMM-Newton MOS with what the use of exposure-weighted solid angle (black points) consistent with observational coverage of the dataset and with direct use of the geometric solid angle (grey points). In order to enable a viable comparison, all fluxes shown here use a 2 mask of the Galactic plane, as provided for the XMM-Newton data.
The discrepancy in solid angle varies between rings, so a full reanalysis is required to derive revised XMM-Newton constraints. Using the same analysis framework as for eROSITA, we compute updated XMM-Newton limits on DM annihilation and decay into $e^{\pm}$, as well as PBH evaporation, shown in Fig. 6. These incorporate propagation uncertainties consistent with those used in the eROSITA analysis. We do not repeat the uncertainty propagation for the DM density distribution or the local DM density as this can be inferred easily from Fig. 4.
For annihilation, the revised XMM-Newton bounds are roughly an order of magnitude weaker than those from eROSITA across the full DM mass range. For decaying DM, XMM-Newton constraints are more competitive at low masses, where FSR dominates, but become nearly two orders of magnitude weaker at higher masses, where IC emission governs the size of the signal. For PBHs, we find that XMM-Newton constraints are consistently about an order of magnitude weaker than those from eROSITA over the full PBH mass range considered. This clarification realigns the role of XMM-Newton in the broader landscape of X-ray DM searches and confirms eROSITA as the current leading X-ray instrument for probing sub-GeV dark matter and evaporating PBHs.
Figure 6: Comparison of the $95\%$ confidence bounds on annihilating (top left panel) and decaying (top right panel) sub-GeV DM and evaporating PBHs (bottom panel) between eROSITA (red) and XMM-Newton (blue) using the exposure-weighted solid angle for the best-fit propagation parameters. In the case of XMM-Newton, we show uncertainties arising from propagation parameters (blue shaded region).
6 Discussion and conclusions
X-ray observations of diffuse Galactic emission offer a powerful avenue to probe new physics scenarios. In particular, sub-GeV DM particles that annihilate or decay into electron-positron pairs populate the interstellar medium with low-energy charged particles. These electrons and positrons interact with Galactic ISRFs via IC scattering and bremsstrahlung, generating diffuse X-ray emission peaking around the keV scale. This makes eROSITA an especially sensitive probe for such signatures, particularly since IC scattering of MeV-scale electrons on the CMB peaks just below 1 keV. Similarly, sub-asteroid mass PBHs emit low-energy $e^{\pm}$ via Hawking evaporation, leading to analogous X-ray signals.
In this work, we used the first all-sky data release from eROSITA (eRASS1) to derive new constraints on sub-GeV DM and asteroid-mass PBHs. We constructed sky maps of the diffuse X-ray background in the 0.3–2.3 keV range and divided the sky into 30 concentric rings centred on the GC, excluding the central $|b|<5^{\circ}$ region to minimise absorption from Galactic gas. These regions were then combined to place 95% confidence level limits.
We find that, even with just six months of data, eROSITA already provides X-ray constraints among the most stringent to date. For decaying DM with mass below $\sim 10$ MeV, our limits surpass those from CMB and diffuse $\gamma$-ray observations. However, in this low-mass regime, signal predictions are subject to larger uncertainties due to charged particle propagation modeling. Looking ahead, eROSITA’s ability to resolve point sources will improve with future data releases, allowing for better subtraction of unresolved backgrounds and potentially yielding significantly stronger constraints. Current diffuse emission is dominated by unresolved sources such as coronal stellar emission and accreting white dwarfs. Improved estimates of this unresolved component could tighten the bounds presented here by orders of magnitude.
We also note that our analysis conservatively excluded the Galactic plane, where DM density is highest. Future work incorporating detailed modeling of X-ray absorption by interstellar gas could improve sensitivity, although CMB constraints will likely remain stronger for annihilating DM.
Finally, we reanalysed previous constraints from XMM-Newton data and identified a discrepancy arising from the treatment of the instrument’s exposure-weighted solid angle. Specifically, an overestimate of the effective sky coverage led to flux values, and therefore constraints, being overstated by approximately four orders of magnitude. After correcting this effect and rederiving the XMM-Newton limits using consistent assumptions, we find that they are weaker than those derived from eROSITA.
Acknowledgements
We thank Joshua Foster for providing useful input regarding the exposure-weighted solid angle for XMM-Newton data analysis. PDL thanks Jari Kajava and Jan-Uwe Ness for their feedback and useful discussions on the topic. SB is supported by the STFC under grant ST/X000753/1. PDL is supported by the Juan de la Cierva JDC2022-048916-I grant, funded by MCIU/AEI/10.13039/501100011033 European Union “NextGenerationEU”/PRTR. The work of PDL is also supported by the grants PID2021-125331NB-I00 and CEX2020-001007-S, wich are both funded “ERDF A way of making Europe” and by MCIN/AEI/10.13039/5011000
11033. PDL also acknowledges the MultiDark Network, ref. RED2022-134411-T. This project used computing resources from the Swedish National Infrastructure for Computing (SNIC) under project Nos. 2021/3-42, 2021/6-326, 2021-1-24 and 2022/3-27 partially funded by the Swedish Research Council through grant no. 2018-05973. DC acknowledges support from a Science and Technology Facilities Council (STFC) Doctoral Training Grant. This project used computing resources from King’s College London CREATE system [65]. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising from this submission.